# Benchmarking 3D Human Pose Estimation Models under Occlusions

### v2

@@ 1 Introduction
Human Pose Estimation (HPE) rom RGB images is a longstanding problem in computer vision field and plays a critical role in numerous real-world applications, including sports[1, 2, 3], healthcare[4, 5, 6, 7], and retail[8]. Specifically, 3D HPE involves estimating the three-dimensional coordinates of key human body points — such as elbows, knees, and wrists — from single images or video sequences. This capability allows systems to accurately interpret and analyze human movement, which is essential for tasks like motion capture in animation[9, 10, 11] and posture assessment in healthcare[5] or sports [3].
@@
Figure 1: BlendMimic3D frames show a subject without (top left) and with occlusion (bottom left). Corresponding 3D ground-truth (top right) and simulated noisy prediction (bottom right) illustrate occlusion impact.
@@
A central challenge in HPE is handling occlusions, where parts of the body are hidden from view—not only due to the subject’s own pose (self-occlusion), but also because of external objects or other people. For instance, in sports such as soccer or basketball, players often block one another, making it difficult for algorithms to reliably estimate keypoint locations[3]. In retail settings, limbs may be occluded by shelves or merchandise, complicating full-body pose detection. Similarly, in clinical gait analysis, assistive devices or medical equipment can obstruct key body parts[4, 6], requiring pose estimation methods to be robust under partial visibility.
@@
Over the years, various models [12, 13, 14, 15, 16, 17, 18, 19, 20] have been developed to address this task, with a significant focus on the method of lifting from 2D to 3D poses. Despite recent advances, model performance continues to decline in the presence of occlusions, highlighting the critical need for improved robustness in 3D HPE systems.
@@
Given this, our focus shifts to practical applications: How effectively can these models perform in real-world scenarios? To this end, we evaluate several state-of-the-art (SOTA) models for which code and pre-trained weights are publicly available [12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
In practical applications, two-stage models first capture a person on camera, detect 2D keypoints, and then lift them to a 3D pose. However, occlusions can lead to inaccuracies in detected 2D poses. As shown in Figure 1, these occlusion scenarios can lead to a decline in the performance of 3D reconstruction. This work focuses on 2D-to-3D lifting models and studies their ability to handle such inaccuracies: Can they effectively predict 3D poses from noisy 2D inputs?
@@
Most existing studies use the Human3.6M [21] dataset, which is a controlled environment with indoor settings and minimal obstructions, except for occasional objects in scenarios such as sitting poses. This setup does not fully represent the challenges of real-world occlusions. We opt for the BlendMimic3D [22] dataset, which includes occlusion labels and synthetic images with occlusions, allowing us to specifically assess model performance under varying occlusion conditions. A challenge we faced was the disparity in keypoint formats across datasets [21, 22, 23]. Thus, we had to adjust the BlendMimic3D format to align it as closely as possible with Human3.6M, thereby facilitating the testing of SOTA models. We then introduce a benchmarking protocol that simulates varying occlusion severities by adding noise into the 2D keypoints labeled as occluded. This allows for systematic comparison of model performance under controlled occlusion conditions. Additionally, we perform per-keypoint sensitivity analyses to identify vulnerabilities in specific body regions.
Our findings reveal several key insights:
  • All models show significant degradation under occlusion, especially on distal keypoints (e.g., feet, wrists).
  • Diffusion-based models, despite their stochastic nature, underperform under occlusion unless equipped with explicit occlusion-aware designs.
  • Certain models (e.g., PoseFormerV2, DTF) show greater resilience due to architectural features such as frequency encoding or dual-view fusion.
  • There exists a critical gap between current benchmarks and real-world challenges, emphasizing the need for more occlusion-aware training and evaluation strategies in 3D HPE.
@@
By analyzing model behavior under realistic occlusion and distribution shifts, this study aims to guide future research toward more robust pose estimation systems.
@@
This paper is structured as follows. Section 2 reviews related work on 3D HPE, with a focus on occlusion handling. Section 3 describes the BlendMimic3D dataset, the selected models, the occlusion simulation protocols, and the evaluation metrics. Section 4 presents benchmarking results under varying occlusion conditions, including a per-keypoint occlusion analysis. Section 5 extends the analysis to the effects of dataset distribution shifts and camera placement. Section 6 discusses the broader challenges and limitations identified in current 3D HPE models. Finally, Section 7 concludes the paper and outlines directions for future research.
@@ 2 Related Work
Deep Learning has significantly advanced 2D Human Pose Estimation (HPE) [24, 25]. However, the same level of progress has not been achieved for 3D HPE, particularly in monocular settings, due to depth ambiguities, insufficient training data, and occlusions [22, 25, 26].
@@
Recent surveys on 3D HPE have reviewed deep learning architectural trends [24], benchmark performances [27], and progress in monocular and multi-view setups [24, 28, 27]. However, these works typically evaluate performance using global accuracy metrics on controlled benchmarks like Human3.6M [21] or MPI-INF-3DHP [29], which do not reflect the complexities of real-world scenarios. Udayan et. al [26] recently conducted a systematic review of monocular 3D HPE, highlighting the gap between academic performance and practical applicability, and emphasizing the need to tackle challenges such as occlusions and depth ambiguity. However, their work is limited to a comparative synthesis of existing literature, they did not conduct new experiments to validate their conclusions.
@@
Sarafinos et al.  [28] tackled the benchmark limitations by introducing a synthetic dataset that allows full control over scene variables and human pose configurations. Their goal was to test model robustness to anthropometric variability, viewpoint changes and action complexity. While this work offered valuable insights into classical methods, it primarily covered pre-deep learning techniques from 2008 to 2015 and included only a limited discussion of modern deep learning-based approaches. Building on this line of inquiry, our work extends the evaluation of occlusion robustness to current state-of-the-art (SOTA) deep learning models, leveraging a dataset specifically designed for fine-grained occlusion analysis. We begin by reviewing the main approaches to monocular 3D HPE, typically divided into direct estimation and 2D-to-3D lifting.
2.1 Direct 3D HPE
Direct methods infer 3D poses directly from RGB images without relying on intermediate 2D keypoint predictions. Li and Chan [30] proposed a dual-branch network leveraging body part correlations, demonstrating the potential of direct regression models. However, their approach lacked the precision and generalization required for complex poses or occluded scenarios.
Pavlakos et al. [31] introduced a voxel-based formulation that predicts 3D heatmaps for each keypoint, addressing depth ambiguity through volumetric likelihoods. However, their method requires post-processing to extract keypoint locations, which introduces non-differentiability and limits spatial precision. Moreover, achieving high resolution in voxel space increases memory and computation costs.
Sárándi et al. [32] tackled the issue of model generalization, arguing that deep learning-based 3D HPE improves when trained under a large amount of data. To address the challenge of merging training data from different datasets, which use inconsistent skeleton formats (i.e., label different anatomical landmarks), they introduce a latent keypoint set as an intermediate representation to unify the data for model training. While this improves training robustness, it does not explicitly enhance occlusion handling.
In contrast, Wang et. al [33] focused specifically on occlusion robustness by proposing Di2Pose, a diffusion-based model that predicts 3D poses as discrete tokens. Rather than relying on large-scale data unification, Di2Pose simulates occlude-and-replace transitions during training to implicitly model occlusions, enabling it to recover missing keypoint information without requiring explicit occlusion labels. Although their model enhance prediction accuracy under occlusions, its inference speed is limited by the iterative nature of diffusion sampling.
Despite these advancements, direct methods still face challenges in resolving depth ambiguities and occlusions due to the absence of intermediate spatial cues and limited contextual reasoning.
@@ 2.1 Direct 3D HPE
Recently, diffusion-based models have emerged as a promising direction for 3D HPE [16, 14, 17, 18, 41], generating and refining multiple 3D hypotheses from 2D observations through a gradual denoising process. There models show potential in addressing challenges such as occlusions and depth ambiguity. Shan et al. [16] introduced D3DP, which generates multiple 3D pose candidates from a single 2D input, using a denoiser conditioned on 2D keypoints and adapted from MixSTE [14]. Choi et al. [17] proposed DiffuPose, which integrates a graph convolutional architecture to capture keypoint topology and guide the denoising process across the graph structure. Xu et al. [18] presented FinePose, a prompt-driven diffusion model that incorporates semantic information, such as body parts, action classes, and motion dynamics, to enable controllable pose generation, which shows promising results in handling complex poses and potential occluded scenarios.
Despite improved accuracy, diffusion models face challenges in computational efficiency due to their iterative nature, requiring multiple forward passes for each prediction. Moreover, their performance remains highly dependent on the quality of 2D inputs. Errors in 2D detection, especially under occlusions, can significantly degrade the final 3D estimation. In this work, we benchmark state-of-the-art models under varying levels of occlusion to assess their robustness and practical viability.
Even in the absence of occlusions, not all keypoints are equally difficult to estimate. In 2D detection, Yu et al. [42] showed that some keypoints are more easily detectable than others due to their scale. However, 3D HPE models typically allocate equal attention to all keypoints. In this work, we show that as the degrees of freedom increase, so does the complexity of 3D estimation for each keypoint, even in non-occluded scenarios.
@@ 2.3 Occlusion Handling in 3D HPE
Our previous work [22] employed graph-based 3D pose refinement block trained on occluded scenarios. While effective in these conditions, this approach does not consistently perform well in non-occluded settings, highlighting the need for occlusion-aware networks. Addressing this, Cheng et al. [43] included a cylinder man model to generate self-occlusion labels. Ghafoor and Mahmood [19] proposed T3D-CNN, an occlusion-guided framework based on temporal dilated convolutions, explicitly modeling missing keypoints via an occlusion guidance matrix. Later, Ghafoor et al. [20] extended this idea with a temporal interpolation-based occlusion guidance mechanism and a transformer-based approach that generates intermediate views, subsequently fused for final 3D pose estimation, a model named DTF. However, both T3D-CNN and DTF [19, 20] rely on the assumption that occlusion labels are known in advance. Their occlusion-guided frameworks depend on an external model to indicate which keypoints are occluded, treating these as missing inputs during prediction. Wang et al. [44] separate features into occlusion-aware and view-dependent components, completing occluded poses in 2D space, potentially leading to inaccurate 3D coordinates. Hardy and Kim [45] reconstruct occluded parts in 3D but do not address all occlusion types. Thus, there is still room for improvement in occlusion-aware models.
@@
Popular 2D HPE datasets include MPII [50], LSP [51], LSP-extended [52] and COCO  [23]. Among these, COCO has become the most widely used benchmark for image-based pose estimation due to its scale and diversity. However, being human-annotated, COCO is also prone to label noise and inconsistencies.
@@
3DPW [56] is the first in-the-wild dataset with 3D ground-truth obtained via wearable IMUs and cameras in real-world environments, including multi-person settings. SURREAL [53] is a synthetic dataset with random combinations of 3D poses, body shapes, textures, and backgrounds. BEDLAM [57], a recent synthetic dataset designed for 3D human pose and shape (HPS) estimation, showed that models trained purely on synthetic data can generalize to real images. Other than these, recent efforts are moving toward more diverse (e.g., occlusion-heavy or multi-person[58]) and sensor-fused datasets.
Despite these recent advances, Human3.6M [21] remains the most widely used benchmark for single-person 3D pose, while COCO [23] dominates in the 2D HPE space. However, each has limitations, COCO’s human curated nature is prone to errors, and Human3.6M lacks occluded and real-world conditions.
Sárándi et al.[59] showed that SOTA 3D HPE models suffer significant performance drops when evaluated under synthetic occlusions. To address this, in our previous work we introduced BlendMimic3D [22], a synthetic dataset designed to include advanced occlusion scenarios. It includes action-specific sequences with keypoint visibility labels per frame, allowing precise evaluation under occlusion.
Most reviews focus on architectural efficiency or dataset diversity but lack a systematic evaluation of model performance under explicit occlusion conditions. In our study, we aim to bridge this gap by analyzing how SOTA 3D HPE models behave under different types and severities of occlusion, using BlendMimic3D dataset [22]. To the best of our knowledge, this is the first benchmark to perform a structured, occlusion-aware evaluation across a diverse set of models.
@@ 3.1 Selected Models
Table 1: Overview of the selected 2D-to-3D human pose estimation models benchmarked in this study. For each model, we summarize its main architectural focus, the availability of pretrained weights, and code references. Occlusion Aware Training (OAT).
@@
VideoPose3D [12, 61] Temporal convolutional network known for its simplicity and competitive performance.
P-STMO [13, 62] Transformer-based model trained with masked keypoint modeling for self-supervised learning.
MixSTE [14, 63] Mixed spatio-temporal transformer that captures fine-grained inter-keypoint motion dynamics.
PoseFormerV2 [15, 64] Frequency-domain transformer designed for long-sequence efficiency and robustness.
D3DP [16, 65] Diffusion-based model using a denoiser adapted from MixSTE.
DiffuPose [17, 66] Diffusion model with a graph convolutional backbone for topology-aware denoising.
FinePose [18, 67] Prompt-driven diffusion model incorporating action class, body part, and kinematic context.
T3D-CNN [19, 68] Occlusion-aware model using temporal convolutions guided by a keypoint-level occlusion mask.
DTF [20, 69] Occlusion-aware transformer-based model with view-fusion and occlusion interpolation.
For completeness, we note that our previously proposed model, a graph convolutional network (GCN) designed to refine 3D poses predicted from occluded 2D inputs [22], was not included in this evaluation. Since the model was trained directly on the BlendMimic3D dataset, using it in the same evaluation would compromise the validity of the benchmark.
@@
T3D-CNN [19](Rand 16) 2022 71.1
DTF-GT [20] (Rand 16) 2024 55.0
This value is higher than in their paper because they use HR-Net [70] as 2D pose detector. We retrained it with CPN detections.
This value is higher than in their paper because they use 2D ground-truth input. We retrained it with CPN detections.
For diffusion-based models, we report results using the Pose-level Aggregation (P-Agg) strategy, where multiple 3D pose hypotheses are averaged to produce a final output. We chose this approach because alternative aggregation strategies, such as JPMA [16], which relies on reprojection accuracy, can become unreliable in the presence of noisy or occluded 2D inputs. Additionally, methods like P-Best and J-Best, which require access to ground-truth 3D poses, are not feasible in real-world applications, where ground-truth 3D poses are unavailable at inference time.
For occlusion-aware models, namely T3D-CNN [19] and DTF [20], we replicate their training protocols using the Rand 16 setting, where 16 keypoints are randomly masked as occluded. These settings correspond to the configurations proposed by the authors in their original papers.
As shown in Table 2, D3DP achieves the best overall performance, followed closely by FinePose and MixSTE. These results highlight the effectiveness of diffusion-based models. Among transformer-based models, MixSTE, P-STMO and PoseFormerV2 all perform competitively, while VideoPose3D, despite its simplicity and earlier design, still offers a solid baseline. However, these numbers represent upper-bound performance, and do not reflect robustness to occlusion, addressed in the following sections.

### v3

@@ 1 Introduction
Human Pose Estimation (HPE) rom RGB images is a longstanding problem in computer vision field and plays a critical role in numerous real-world applications, including sports[46, 3, 55], healthcare[6, 29, 30, 2], and retail[25]. Specifically, 3D HPE involves estimating the three-dimensional coordinates of key human body points — such as elbows, knees, and wrists — from single images or videos. This capability allows systems to accurately interpret and analyze human movement, which is essential for tasks like motion capture in animation[20, 51, 42] and posture assessment in healthcare[29] or sports [55].
@@
Figure 1: BlendMimic3D frames show a subject without (top left) and with occlusion (bottom left). Corresponding 3D ground-truth (top right) and noisy prediction (bottom right) illustrate occlusion impact.
@@
A central challenge in HPE is handling occlusions, where parts of the body are hidden from view—not only due to the subject’s own pose (self-occlusion), but also because of external objects or other people. For instance, in sports such as soccer or basketball, players often block one another, making it difficult for algorithms to reliably estimate keypoint locations[55]. In retail settings, limbs may be occluded by shelves or merchandise, preventing full-body pose detection. Similarly, in clinical gait analysis, assistive devices or medical equipment can obstruct key body parts[6, 30], requiring pose estimation methods to be robust under partial visibility.
@@
Over the years, various models [32, 38, 57, 59, 39, 9, 53, 12, 11, 36, 16, 33, 14, 5, 24] have been developed to address 3D HPE, with a significant focus on the method of lifting from 2D to 3D poses. Despite recent advances, model performance continues to decline in the presence of occlusions, highlighting the critical need for improved robustness in 3D HPE systems.
@@
Given this, our focus is on understanding how well current 3D HPE models handle occlusions, which are inherently present in most real-world settings. To this end, we evaluate several state-of-the-art (SOTA) models for which code and pre-trained weights are publicly available [32, 38, 57, 59, 39, 9, 53, 12, 11].
@@
In 3D HPE, two-stage models first capture a person on camera, detect 2D keypoints, and then lift them to a 3D pose. However, occlusions can lead to inaccuracies in detected 2D poses. As shown in Figure 1, these occlusion scenarios can lead to a decline in the performance of 3D reconstruction. This work focuses on 2D-to-3D lifting models and studies their ability to handle such inaccuracies: Can they effectively predict 3D poses from noisy 2D inputs?
@@
Most existing studies use the Human3.6M [15] dataset, which is a controlled environment with indoor settings and minimal obstructions, except for occasional objects in scenarios such as sitting poses. This setup does not fully represent the challenges of real-world occlusions. We opt for the BlendMimic3D [24] dataset, which includes occlusion labels and synthetic images with occlusions, To the best of our knowledge, BlendMimic3D is the only 3D HPE dataset that provides explicit, frame-level and keypoint-level occlusion labels. These occlusion annotations are central to this work, as they allow us to simulate occlusion noise and analyze how SOTA models respond to it. Thus, we propose a benchmarking protocol for systematic comparison of model performance under controlled occlusion conditions. Additionally, we perform per-keypoint sensitivity analyses to identify vulnerabilities in specific body regions.
To summarize, our work makes the following contributions:
  • We introduce a comprehensive benchmark of SOTA 3D HPE models under realistic occlusion conditions, moving beyond clean-input and randomly masked evaluation settings.
  • Through a controlled occlusion simulation based on 2D detector behavior, we reveal differences in robustness across convolutional, transformer-based, graph-based, and diffusion-based lifting architectures.
  • We propose two evaluation protocols, supporting both global and per-keypoint performance analyses, to assess model robustness to occluded 2D keypoints.
  • Our results show consistent vulnerability patterns across models, with distal keypoints (e.g., wrists and feet) among the most error-prone.
  • Finally, we show that recent diffusion-based and occlusion-aware approaches, while effective in controlled settings, do not inherently confer robustness to realistic occlusion patterns, highlighting open challenges for deploying 3D HPE systems in unconstrained environments.
@@
This paper raises critical questions about the applicability of 3D HPE advances in real-world settings and the adequacy of current benchmarks. By analyzing model behavior under realistic occlusion and distribution shifts, this study aims to guide future research toward more robust pose estimation systems.
@@
This paper is structured as follows. Section 2 reviews related work on 3D HPE, with a focus on occlusion handling. Section 3 describes the BlendMimic3D dataset, the selected models, the occlusion simulation protocols, and the evaluation metrics. Section 4 presents benchmarking results under varying occlusion conditions, including a per-keypoint occlusion analysis. Section 5 extends the analysis to the effects of dataset distribution shifts and camera placement. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 2 Related Work
Deep Learning has significantly advanced 2D Human Pose Estimation (HPE) [61, 54]. Although 3D HPE has also benefited from these advantages, it remains more challenging, particularly in monocular settings, due depth ambiguities, limited training data, and occlusions [24, 54, 43].
@@
Recent surveys on 3D HPE have reviewed deep learning architectural trends [61], benchmark performances [47], and progress in monocular and multi-view setups [61, 35, 47]. However, these works typically evaluate performance using global accuracy metrics on controlled benchmarks like Human3.6M [15] or MPI-INF-3DHP [28], which do not reflect the complexities of real-world. Udayan et. al [43] recently conducted a systematic review of monocular 3D HPE, highlighting the gap between academic performance and practical applicability, and emphasizing the need to tackle challenges such as occlusions and depth ambiguity. However, their work focuses on synthesizing existing literature and does not include new experimental validation of the highlighted challenges.
@@
Sarafinos et al.  [35] tackled the benchmark limitations by introducing a synthetic dataset that allows full control over scene variables and human pose configurations. Their goal was to test model robustness to anthropometric variability, viewpoint changes and action complexity. While this work offered valuable insights into classical methods, it primarily covered pre-deep learning techniques from 2008 to 2015 and included only a limited discussion of modern deep learning-based approaches. Building on this line of research, our work extends the evaluation of occlusion robustness to current state-of-the-art (SOTA) deep learning models, leveraging a dataset specifically designed for fine-grained occlusion analysis.
3D HPE models are typically divided into direct estimation and 2D-to-3D lifting. Some works have explored direct 3D HPE, which estimates 3D poses directly from images without relying on intermediate 2D detections [22, 31, 36, 49]. However, direct methods face challenges in resolving depth ambiguities and occlusions due to the absence of intermediate spatial cues and limited contextual reasoning.
In this work, we focus on lifting approaches, which remain widely used due to their ability to leverage existing 2D pose detectors as input. In the following sections, we review the main approaches to monocular 2D-to-3D HPE models.
@@
Recently, diffusion-based models have emerged as a promising direction for 3D HPE [39, 57, 9, 53, 21], generating and refining multiple 3D hypotheses from 2D observations through a gradual denoising process. These models show potential in addressing challenges such as occlusions and depth ambiguity [39, 9, 53]. Shan et al. [39] introduced D3DP, which generates multiple 3D pose candidates from a single 2D input, using a denoiser conditioned on 2D keypoints and adapted from MixSTE [57]. Choi et al. [9] proposed DiffuPose, which integrates a graph convolutional architecture to capture keypoint topology and guide the denoising process across the graph structure. Xu et al. [53] presented FinePose, a prompt-driven diffusion model that incorporates semantic information, such as body parts, action classes, and motion dynamics, to enable controllable pose generation, which shows promising results in handling complex poses.
Despite improved accuracy, diffusion models face challenges in computational efficiency due to their iterative nature, requiring multiple forward passes for each prediction. Their performance is also highly dependent on the quality of 2D inputs, making them vulnerable to occlusions. Furthermore, even in the absence of occlusions, not all keypoints are equally difficult to estimate. In 2D detection, Yu et al. [56] showed that some keypoints are more easily detectable than others due to their scale. However, 3D HPE models typically allocate equal attention to all keypoints.
@@ 2.2 Occlusion Handling in 3D HPE
Our previous work [24] employed graph-based 3D pose refinement block trained on occluded scenarios. While effective in these conditions, this approach does not consistently perform well in non-occluded settings, highlighting the need for occlusion-aware networks. Addressing this, Cheng et al. [8] included a cylinder man model to generate self-occlusion labels. Ghafoor and Mahmood [12] proposed T3D-CNN, an occlusion-guided framework based on temporal dilated convolutions, explicitly modeling missing keypoints via an occlusion guidance matrix. Later, Ghafoor et al. [11] extended this idea with a temporal interpolation-based occlusion guidance mechanism and a transformer-based approach that generates intermediate views, subsequently fused for the final 3D pose estimation, a model named DTF. However, both T3D-CNN and DTF [12, 11] rely on the assumption that occlusion labels are known in advance. Their occlusion-guided frameworks depend on an external model to indicate which keypoints are occluded, treating these as missing inputs during prediction. Wang et al. [48] separate features into occlusion-aware and view-dependent components, completing occluded poses in 2D space, potentially leading to inaccurate 3D coordinates. Hardy and Kim [13] reconstruct occluded parts in 3D but do not address all occlusion types. Thus, there is still room for improvement in occlusion-aware models.
@@
Popular 2D HPE datasets include MPII [1], LSP [17], LSP-extended [18] and COCO  [23]. Among these, COCO has become the most widely used benchmark for image-based pose estimation for its scale and diversity. However, being human-annotated, COCO is also prone to label noise and inconsistencies.
@@
3DPW [45] is the first dataset with 3D ground-truth obtained via wearable IMUs and cameras in real-world environments, including multi-person settings. SURREAL [44] is a synthetic dataset with random combinations of 3D poses, body shapes, textures, and backgrounds. BEDLAM [4], a recent synthetic dataset designed for 3D human pose and shape (HPS) estimation, showed that models trained purely on synthetic data can generalize to real images. Other than these, recent efforts are moving toward more diverse (e.g., occlusion-heavy or multi-person[10]) and sensor-fused datasets.
Despite recent advances, Human3.6M [15] remains the main benchmark for single-person 3D pose, while COCO [23] dominates in 2D HPE. However, each has limitations, COCO’s human curated nature is prone to errors, and Human3.6M lacks occluded and real-world conditions.
Sárándi et al.[37] showed that SOTA 3D HPE models suffer significant performance drops when evaluated under synthetic occlusions. To address this, in our previous work we introduced BlendMimic3D [24], a synthetic dataset designed to include realistic occlusion scenarios. It includes action-specific sequences with keypoint visibility labels per frame, allowing precise evaluation under occlusion. Building on this resource, the present study leverages BlendMimic3D to analyze the behavior of SOTA 3D HPE models under different types and severities of occlusion. To the best of our knowledge, this is the first benchmark to perform a structured, occlusion-aware evaluation across a diverse set of models.
@@ 3.1 Selected Models
Table 1: Overview of the selected 2D-to-3D human pose estimation models benchmarked in this study. For each model, we summarize its main architectural focus, if it involves an occlusion aware training (OAT), and the availability of pretrained weights.
@@
VideoPose3D [32] Temporal convolutional network known for its simplicity and competitive performance.
P-STMO [38] Transformer-based model trained with masked keypoint modeling for self-supervised learning.
MixSTE [57] Mixed spatio-temporal transformer that captures fine-grained inter-keypoint motion dynamics.
PoseFormerV2 [59] Frequency-domain transformer designed for long-sequence efficiency and robustness.
D3DP [39] Diffusion-based model using a denoiser adapted from MixSTE.
DiffuPose [9] Diffusion model with a graph convolutional backbone for topology-aware denoising.
FinePose [53] Prompt-driven diffusion model incorporating action class, body part, and kinematic context.
T3D-CNN [12] Occlusion-aware model using temporal convolutions guided by a keypoint-level occlusion mask.
DTF [11] Occlusion-aware transformer-based model with view-fusion and occlusion interpolation.
@@
T3D-CNN [12] (None) 2022 56.6
DTF-GT [11] (Rand 16) 2024 53.5
This value is higher than in their paper because they use HR-Net [41] as
2D pose detector. We retrained it with CPN detections.
This value is higher than in their paper because they use 2D ground-truth
input. We retrained it with CPN detections.
For diffusion-based models, we report results using the Pose-level Aggregation (P-Agg) strategy, where multiple 3D pose hypotheses are averaged to produce a final output. We chose this approach because alternative aggregation strategies, such as Joint-wise reProjection-based Multi-hypothesis Aggregation (JPMA) [39], which relies on reprojection accuracy, can become unreliable in the presence of noisy or occluded 2D inputs. Additionally, methods like P-Best and J-Best, which require access to ground-truth 3D poses, are not feasible in real-world applications, where ground-truth 3D poses are unavailable at inference time.
For occlusion-aware models, namely T3D-CNN [12] and DTF [11], we replicate the training protocols proposed by the original authors. For DTF, we report results using the Rand 16 setting, where 16 keypoints are randomly masked as occluded. In contrast, for T3D-CNN we select the None configuration, in which the occlusion-guidance channels are removed, as it yields the most stable performance among the evaluated T3D-CNN variants. A more detailed analysis of T3D-CNN under different masking strategies is provided later in the paper.
As shown in Table 2, D3DP achieves the best overall performance, followed closely by FinePose and MixSTE. These results highlight the effectiveness of diffusion-based models in occlusion-free settings. Among transformer-based models, MixSTE, P-STMO and PoseFormerV2 all perform competitively, while VideoPose3D, despite its simplicity and earlier design, still offers a solid baseline. However, these numbers represent upper-bound performance, and do not reflect robustness to occlusion, addressed in the following sections.